The table shows the population of elk in a state forest growing exponentially. Determine the equation that expresses the number of elk, y, present at any time, t.
Time (yrs.)------------Population
0-------------------------30
1-------------------------42
3------------------------82.32
6------------------------161.35

Answer by Theo(13342)

30 * 1.4^1 = 42
30 * 1.4^3 = 82.32
30 * 1.4^5 = 161.3472

i suspect that your 6 years should be 5 years.
it works with 5.
it doesn't work with 6.
if it's exponential, the rate should be the same for all.
the equation, as far as i know, should be y = 30 * 1.4 ^ x
here's my graph of the equation.



the exponential formula is y = a * b ^ x
in this problem, a = 30 and b = 1.4